Consider the polynomial function 𝑃(π‘₯) = 2π‘₯^3 βˆ’ π‘šπ‘₯^2 + π‘₯ βˆ’ 5π‘š. The remainder when P(x) is divided by (π‘₯ βˆ’ 2) is four times the remainder from dividing P(x) by (π‘₯ + 1). Determine m algebraically and show all your work.Β 


1
Expert's answer
2021-11-24T17:28:30-0500

"\text{Since x-2 and x+1 divides P(x), we have that at x = 2 and x = -1, therefore}\\
2(2)^3-m(2)^2+2-5m = R(x)\\
2(-1)^3 - m (-1)^2+ 2 -5m=R(x)\\
\text{Therefore, we have }\\
18-9m= -12-24m\\
15m=-30\\
\implies m = -2"